OpenStudy (marigirl):

Integration question. The question is to find the integral of -cosec6xcis6x Plz help! I thought to first integrate -cosec6x Which is 1/sin (6x).....a little stuck

OpenStudy (nincompoop):

can you rewrite that properly

OpenStudy (marigirl):

Integrate -cosec6xcot6x

OpenStudy (anonymous):

@nincompoop can you plz help me in a ques?

OpenStudy (nincompoop):

\(\huge \int -csc (6x) cot(6x) dx \) like this?

OpenStudy (marigirl):

Does cse mean cosec? If so then yes

OpenStudy (anonymous):

Ok so rewrite it as -1/sin(6x) * cos(6x)/sin(6x) then you get -cos(6x)/sin^2(6x) Then find the derivative of sin(6x) then your finished and use the formula integ of u^ndu=u^n+1/n+1 +c

OpenStudy (nincompoop):

\(\huge -\frac{1}{sin} = -csc \)

OpenStudy (nincompoop):

\(\huge \frac{1}{tan} = cot \)

OpenStudy (marigirl):

Does cse mean cosec? If so then yes

OpenStudy (nincompoop):

yes

OpenStudy (nincompoop):

u-sub

OpenStudy (anonymous):

Look thats it

OpenStudy (nincompoop):

u = 6x du = 6 dx

OpenStudy (anonymous):

Did you understand it?

OpenStudy (nincompoop):

v = sin(u) dv = cos(u) du

OpenStudy (nincompoop):

those are you keys

OpenStudy (anonymous):

Your description and point is surely very feeble my friend so you're saying the Bible is demonic?

OpenStudy (anonymous):

In asking you a question that relates to what you've said hence fabricator

OpenStudy (nincompoop):

you really need to read the whole thing before you ruin your mind with fallacy

OpenStudy (anonymous):

No need since i spotted an error just from the very beginning which discourages me to continue this nonsense. Have a great day